OpenStudy (lilmike234):

what is the radius of a circle with the equation x^2+y^2+2x+4y-9=0

Nnesha (nnesha):

standard form equation of the circle \[\huge\rm (x-h)^2+(y-k)^2=r^2\] where (h,k) is the center point and r =radius

Nnesha (nnesha):

so move the -9 to the right side

OpenStudy (jdoe0001):

do you know what a "perfect square trinomial" is? sometimes also just called a "perfect square"

OpenStudy (lilmike234):

Yeah I heard of it.

OpenStudy (zzr0ck3r):

\(x^2+y^2+2x+4y-9=0\\ x^2+y^2+2x+4y=9\\ x^2+2x+y^2+4y=9\\ (x+1)^2-1+y^2+4y=9\\ (x+1)^2+y^2+4y=10\) Now do something similar with the \(y^2+4y\).

OpenStudy (lilmike234):

So I would get 3 for my answer correct?

OpenStudy (zzr0ck3r):

Now look at what @Nnesha said

OpenStudy (lilmike234):

So my answer would be 6?

OpenStudy (zzr0ck3r):

\(x^2+y^2+2x+4y-9=0\\ x^2+y^2+2x+4y=9\\ x^2+2x+y^2+4y=9\\ (x+1)^2-1+y^2+4y=9\\ (x+1)^2+y^2+4y=10\\ (x+1)^2+(y+2)^2-4=10\\ (x+1)^2+(y+2)^2=14\\ (x+1)^2+(y+2)^2=(\sqrt{14})^2\)

OpenStudy (lilmike234):

So my answer should be 3.74 correct ?

OpenStudy (zzr0ck3r):

how did you get that?

OpenStudy (lilmike234):

I square rooted 14

OpenStudy (zzr0ck3r):

right